WAR (We Are Rich) has been in business since 1986. WAR is an accrual method sole proprietorship that deals in the manufacturing and wholesaling of various types of golf equipment. Hack & Hack CPAs has filed accurate tax returns for WAR’s owner since WAR opened its doors. The managing partner of Hack & Hack (Jack) has gotten along very well with the owner of WAR—Mr. Someday Woods (single). However, in early 2019, Jack Hack and Someday Woods played a round of golf and Jack, for the first time ever, actually beat Mr. Woods. Mr. Woods was so upset that he fired Hack & Hack and has hired you to compute his 2019 taxable income. Mr. Woods was able to provide you with the following information from prior tax returns. The taxable income numbers reflect the results from all of Mr. Wood’s activities except for the items separately stated. You will need to consider how to handle the separately stated items for tax purposes. Also, note that the 2014–2018 numbers do not reflect capital loss carryovers.

2014 2015 2016 2017 2018
Ordinary taxable income $ 4,000 $ 2,000 $ 94,000 $ 170,000 $ 250,000
Other items not included in ordinary taxable income:
Net gain (loss) on disposition of §1231 assets $ 3,000 10,000 $ (6,000 )
Net long-term capital gain (loss) on disposition of capital assets $ (15,000 ) $ 1,000 $ (7,000 ) $ (7,000 )

In 2019, Mr. Woods had taxable income in the amount of $480,000 before considering the following events and transactions that transpired in 2019:

  1. On January 1, 2019, WAR purchased a plot of land for $100,000 with the intention of creating a driving range where patrons could test their new golf equipment. WAR never got around to building the driving range; instead, WAR sold the land on October 1, 2019, for $40,000.
  2. On August 17, 2019, WAR sold its golf testing machine, “Iron Byron” and replaced it with a new machine “Iron Tiger.” “Iron Byron” was purchased and installed for a total cost of $22,000 on February 5, 2015. At the time of sale, “Iron Byron” had an adjusted tax basis of $4,000. WAR sold “Iron Byron” for $25,000.
  3. In the months October through December 2019, WAR sold various assets to come up with the funds necessary to invest in WAR’s latest and greatest invention—the three-dimple golf ball. Data on these assets are provided below:
Asset Placed in Service (or purchased) Sold Initial Basis Accumulated Depreciation Selling Price
Someday’s black leather sofa
(used in office)
4/4/18 10/16/19 $ 3,000 $ 540 $ 2,900
Someday’s office chair 3/1/17 11/8/19 8,000 3,000 4,000
Marketable securities 2/1/16 12/1/19 12,000 0 20,000
Land held for investment 7/1/18 11/29/19 45,000 0 48,000
Other investment property 11/30/17 10/15/19 10,000 0 8,000

d. Finally, on May 7, 2019, WAR decided to sell the building where it tested its plutonium shaft, lignite head drivers. WAR purchased the building on January 5, 2007, for $190,000 ($170,000 for the building, $20,000 for the land). At the time of the sale, the accumulated depreciation on the building was $50,000. WAR sold the building (with the land) for $300,000. The fair market value of the land at the time of sale was $45,000. (Do not round intermediate computations. Round your final answers to the nearest whole dollar amount. Loss amounts should be indicated by a minus sign.)

Compute Mr. Woods’s taxable income after taking into account the transactions described above

Compute Mr. Woods's tax liability for the year. (Ignore any net investment income tax for the year and assume the 20 percent qualified business income deduction is included in taxable income before these transactions.) Use Tax rate schedules
Answer #1

part 1

Item

Amount

Notes

Character

TI before adjustments

480000

Ordinary

Loss from sale of land held < 1 year

(60000)

A

Ordinary

Sec. 1245 recapture

18440

B

Ordinary

Sec. 1250 recapture

0

C

Sec. 1231 gain recaptured as ordinary (5-yr lookback rule)

6000

D

Ordinary

Net Sec. 1231 gain (after application of look back rule)

156000

D

Capital

Net long-term capital gain from investments (before capital loss carry forwards)

9000

E

Capital

Capital loss carry- forwards

(3000)

F

Capital

TI after adjustments

$606440

  1. Cash received $100000 with a $40,000 basis giving a loss of $60000. As property is not held for more than one year, it is not categorized as section 1231 property and thus loss is ordinary loss.
  2. Recaptured gain according to Sec. 1245 refers to gain on the sale of an asset that results because lowered basis due to depreciation. Gain = 25000-4000 = $21000. Recaptured gain as ordinary income under Sec. 1245 = $18000 (lowered tax basis due to depreciation). Remaining $3000 is Sec. 1231 gain. Sofa is a Sec. 1231 property and gain on its disposal = $440 (2900-(3000-540)). Recaptured gain as ordinary income under Sec. 1245 = $440 (lowered tax basis due to depreciation). Therefore total Sec. 1245 gain = 18000+440 = $18440
  3. Section 1250 recapture is applicable to real property that is depreciated as per the accelerated method of depreciation. amount of recapture is the difference of accelerated depreciation amount and depreciation amount using straight line method. Selling price of building = selling price for the land and the building - fair market value of the land = 300000-45000 = $255000. Gain on sale = 255000-135000 = $120000. Thus there is no recapture under Section 1250.
  4. net Section 1231 gain = 3000+135000+(45000-20000) -1000 = 162000

As 5 yr. lookback rule applies to net gain, the loss of $6000 will be recognized. Thus net gains of section 1231 of $156000 will be treated as long-term capital gains.

  1. long-term capital gain on the sale of securities = 20000-12000 = 8000

long-term capital gain of on the land held for investment = 48000-45000 = 3000

long-term capital loss on the sale of other investment property = -2000

thus, net long-term capital gain = $9000

  1. long-term capital loss= 3000

Part 2

Ordinary income

Taxable income before transactions

480,000

Ordinary loss – land (40000-100000)

(60000)

Recapture (22000-4000)+(2900-(3000-540))

18440

Ordinary income from lookback

6000

Ordinary income

444440

LTCG at 25%

§1250 Gain – Bldg

50000

Ordinary income from lookback

6000

44000

LTCL carryforward

(7000)

Net LTCG at 25%

37000

LTCG at 0/15/20%

Gains:

Iron Byron (25000-22000)

3000

Building (255000-120000) = 135000 of which 50000 Unrecaptured §1250 gain

85000

Land (45000-20000)

25000

Marketable Securities (20000-12000)

8000

Land – Investment (48000-45000)

3000

Losses:

Chair (4000-5000)

(1000)

Investment Property (8000-10000)

(2000)

Net LTCG

121000

Tax Liability:

Tax on ordinary income

130747 (46628+(35%*(444440-204100)))

Tax on 25% Gain

9250 (37000*25%)

Tax on 0/15/20% Gain (taxed at 20%)

24200 (121000*20%)

Total Tax liability

164197